# Steering Embedding Models with Geometric Rotation: Mapping Semantic Relationships Across Languages and Models

### v1

@@ 6.1 Cross-Language Transfer Comparison
This section discusses the comparison of models trained in one of the seven languages and tested on all seven. The results of this section demonstrate RISE multilingual performance computed by three embedding models. See Appendix B for comprehensive results across all phenomena for each model.
Negation emerges as the most robust discourse-level, semantic transformation, achieving the highest mean rotor alignment score (0.788) across all model-language combinations with performance ranging from 0.686 to 0.918. Figure 1 demonstrates RISE performance on negation for each model. RISE transformations for negation are most geometrically consistent in text-embedding-3-large. Negation’s strong performance indicates that generalizable discourse-level, semantic changes are captured by RISE and best applied cross-lingually in text-embedding-3-large.
@@
The cross-language analysis presented in Appendix B supports our hypothesis that discourse-level semantic transformations correspond to intrinsic geometric operations on the embedding manifold, rather than fixed directions derived from Euclidean computations. The variation across models, preservation of linguistic relationships across languages, and transformation patterns indicate that RISE successfully identifies sematic transformation on the embedding manifold. The limitations and future work are discussed further on.
@@ 6.2 Cross-Model Transfer Comparison
To evaluate RISE prototypes’ robustness to transfer across different embedding architectures, we conducted cross-model mapping experiments using the method developed by Morris et al. (2020). This approach learns statistical mappings between embedding spaces through principal component analysis and distributional alignment, enabling transfer of learned RISE prototypes from one model to another. We specifically examined transfer from text-embedding-3-large (3072-dimensional) to bge-m3 (1024-dimensional), demonstrating cross-model semantic transfer across different dimensionalities and training objectives. For each language pair and phenomenon, we learn RISE prototypes in text-embedding-3-large using 80% of the data, map these prototypes and $e_{1}$ to bge-m3 space, and evaluate performance on native bge-m3 embeddings using the remaining 20%. Figure 3 demonstrates comprehensive cross-model and cross-language transfer results.
@@ 6.3 English Task-Based Comparison
Our main investigation is how well RISE peforms in in multi-lingual settings. However there are limited external datasets for evaluating the performance discourse-level, semantics transformation tasks. Due to the limited resources, we had to select the most related datasets, BLiMP and SICK. BLiMP is LLM evaluation paired sentence dataset for major grammatical phenomena in English, and SICK is a dataset with paired sentences with entailment, contradiction, and neutral labels.
Table 1 summarizes RISE performance across the three datasets. The results confirm that all models achieve strong performance, with particular strengths varying by dataset: mBERT excels on grammatical tasks (BLiMP) and contradiction detection (SICK), while bge-m3 shows the most consistent performance across synthetic multilingual data. The dramatic performance gap between BLIMP (>>0.92) and SICK (0.62-0.74) suggests that RISE rotations might be capturing something more specific than general semantic transformations.
The high BLIMP performance indicates RISE excels at preserving grammatical/syntactic structure, while the moderate SICK performance suggests these same rotations don’t preserve semantic relatedness as well. These results show that benchmark choice dramatically affects relative model ranking. Instead, robustness depends on whether the task prioritizes cross-lingual consistency (favoring bge-m3) or raw performance on specific phenomena (favoring text-embedding-3-large for negation, mBERT for grammatical tasks).
Table 1: RISE Validation: Performance Across Three Validation Datasets. The performance is measured with the rotor alignment score between RISE-steered embeddings and target embeddings where bold values indicate best performance per dataset.
@@
7 Discussion
@@
Our findings demonstrate that meaningful semantic operations can be recovered as geometric transformations in modern language model representations. RISE successfully identifies consistent geometric structure for discourse-level semantic changes, primarily for text-embedding-3-large and negation in multilingual settings. These findings provide positive results for extending the interpretable geometry of early word embeddings to contemporary embeddings in high dimensional spaces.
@@
Evaluation benchmarks (Table 1) reveal task-dependent effectiveness. RISE achieves near-perfect performance on syntactic acceptability (BLIMP: 0.93-0.96) but only moderate performance on semantic similarity (SICK: 0.62-0.74), suggesting better alignment with grammatical rather than semantic transformations. Section 6.1 shows that negation and conditionality are the most generalizable discourse-level, semantic changes captured by RISE and best applied cross-lingually in text-embedding-3-large. Together these results support that RISE is most successful at identifying semantic transformation with distinct grammatical factors, but more work is needed to justify semantic transformations in multilingual models are universal geometric operations.
8 Limitations and Future Work
@@
First, our analysis focuses on three specific linguistic transformation types. Future work should expand to additional semantic and pragmatic phenomena to test the generality of geometric consistency principles. Second, while our experiments used three diverse embedding models (text-embedding-3-large, bge-m3, and mBERT), validation across additional architectures would strengthen claims about the universality of geometric semantic structure. Third, while our language selection spans diverse typological features, the reliance on GPT-4.5 for data generation may introduce subtle biases toward English-centric conceptualizations of semantic phenomena. Future work should incorporate more diverse data sources and validation by native speakers.
9 Broader Implications for Multilingual Model Interpretability
The cross-linguistic transfer results connect to broader research on multilingual representation learning (Feng et al., 2022; Chang et al., 2022). Our findings provide support that modern multilingual embedding models share similar geometric structures for semantic operations with distinct grammatical factors (e.g. negation and conditionality), but more work is needed to understand how complex semantic transformation (e.g. politeness) are represented by theories of universal semantic representation. These results have implications for language model interpretability. The demonstrated ability to learn geometric transformations for semantic operations suggests that: (1) Despite their complexity, modern language models maintain interpretable geometric structure for certain types of semantic relationships. (2) The cross-linguistic consistency of geometric transformations suggests that multilingual models learn universal semantic representations that transcend language-specific surface features. (3) The ability to predict semantic transformations geometrically opens possibilities for more principled control of model behavior through targeted geometric interventions.
@@ 10 Conclusion
The ability to learn geometric transformations for discourse changes relates to work on controllable text generation and steering vectors (Turner et al., 2023; Li et al., 2023). Our rotor-based approach, RISE, provides a geometric framework for understanding and implementing semantic control in language models. This work investigated whether discourse-level semantic transformations in multilingual embedding spaces correspond to intrinsic geometric operations, specifically rotations identified through the RISE method. Our comprehensive evaluation across multiple models, languages, and validation tasks reveals a more complex reality than initially hypothesized. This work demonstrates that modern language model representations maintain interpretable geometric structure for some semantic transformations, extending the promise of geometric semantics from early word embeddings to contemporary transformer models. We show that:
@@
    Semantic transformations exhibit consistent geometric structure, with consistency varying systematically based on context-independence.
@@
    Cross-linguistic transfer of geometric transformations achieves consistent performance, supporting theories of universal semantic representation in multilingual models.
  3. 3.
    The RISE method provides a framework for identifying semantically meaningful geometric structure in high-dimensional representation spaces.
As language models continue to evolve, understanding these geometric foundations will be crucial for developing more interpretable and controllable AI systems. By revealing transferable geometric structure in semantic transformations (e.g. negation and conditionality), this work opens new possibilities for understanding and controlling language model behavior through geometric interventions. Our work establishes geometric steering as a principled approach to cross-lingual semantic control, achieving 77%-95% cross-language transfer effectiveness across typologically diverse languages. By developing RISE, we demonstrate that interpretable structure exists for some grammatically distinct semantic transformations, providing a tools for understanding and controlling how these systems encode semantic knowledge. While RISE remains valuable for analyzing model-specific semantic structures, claims about universal geometric operations require substantial qualification. Our cross-model transfer experiments expose an English-centric bias, with English achieving 20% higher transfer scores than languages like Zulu. This English-centric bias persists across all semantic transformations, indicating that current multilingual models encode geometric structures that prioritize English. Future work should focus on developing more equitable multilingual representations and investigating which language-specific geometric structures are an inherent feature of the models.
@@ Appendix A Mathematical Properties of RISE
Roadmap.
This appendix has two parts. First, we state geometry preliminaries on the unit sphere, including explicit exponential and logarithmic map formulas (Lemma 1). Second, we analyze sequential RISE edits: Theorem A.1 proves that RISE transformations commute up to second order in prototype magnitudes, and Proposition A.1 shows that each RISE update scales linearly in the embedding dimension $d$. Together these results provide a rigorous foundation for RISE’s geometric behavior and computational efficiency.
Relevance.
@@ Appendix B Cross-Language Transfer Analysis and Results
To test whether geometric transformations generalize across languages, we conducted comprehensive cross-language transfer experiments. This section reports detailed results across 3 models and 3 semantic phenomena, analyzing both quantitative performance and geometric properties of learned transformations.
@@ B.1 Cross-Language Transfer Performance
The above heatmaps demonstrate comprehensive cross-language transfer results across our three models. Training rotor prototypes on one language and evaluating on others reveals remarkable cross-linguistic performance, particularly for negation and conditionality. Most language pairs show transfer scores above 0.70, with negation achieving particularly strong off-diagonal performance (most scores >> 0.80).
@@ B.2 Geometric Analysis of Cross-Language Centroids
Analysis of the learned centroids reveals additional insights into the geometric structure of semantic transformations. For each phenomenon, we computed ”ideal” transformation vectors by averaging canonicalized transformed embeddings across languages.
@@ B.3 Quantitative Cross-Language Analysis
Model Phenomenon All Transfers Monolingual Cross-Lang Ratio OpenAI (3072d) Conditionality 13.6× ± 0.7 14.5× 13.5× 0.93 Negation 19.7× ± 1.2 20.6× 19.6× 0.95 Politeness 23.1× ± 1.9 25.3× 22.8× 0.90 BGE-M3 (1024d) Conditionality 13.9× ± 0.6 14.5× 13.8× 0.95 Negation 18.5× ± 0.8 19.3× 18.4× 0.95 Politeness 25.2× ± 1.2 26.4× 25.1× 0.95 mBERT (768d) Conditionality 12.8× ± 1.3 15.0× 12.5× 0.83 Negation 18.0× ± 1.8 20.9× 17.5× 0.84 Politeness 20.8× ± 3.9 26.1× 20.0× 0.77
@@
Model Dims Validation Avg Cross-Lang Avg Random Adv OpenAI text-embedding-3-large 3072 0.774 19.0× 6.3× BGE-M3 1024 0.790 19.8× 11.7× mBERT 768 0.802 16.9× 11.9×
@@
Tables 2 and 3 provide comprehensive quantitative analysis of cross-language transfer performance. Notably, all models maintain strong cross-language performance (77%–95% of monolingual performance), with bge-m3 showing the most consistent cross-language effectiveness across all phenomena.
Appendix C Prompt Templates
@@
C.1 Neutral Sentence Generation
@@ C.4 Conditionality
Appendix D Data Generation Methodology
D.1 Diversity Controls
To guard against artefacts that might arise from narrow lexical or topical coverage we apply several sampling diversifiers. (i) Each neutral sentence prompt draws its situation description from a randomly chosen context category and exemplar, yielding a wide topical spread before any transformation is applied. (ii) Within every language we shuffle sentence–feature assignments so that no specific lexical field correlates with a particular transformation subtype. (iii) For each transformation we uniformly sample property values (e.g., negation particle, politeness strategy) per language and sentence, guaranteeing that every combination of language and subtype appears the same number of times. (iv) After generation we remove near-duplicates and enforce a 5–25 token length window, which empirically yields a near-uniform length distribution. Together these steps ensure that our corpus varies in topic, syntax, and lexical choice while remaining balanced across languages and transformation subtypes. These controls ensure that observed geometric patterns reflect semantic properties rather than artifacts of lexical choice or sentence structure.

### v2

@@ 6.1 Cross-Language Transfer Comparison
This section discusses the comparison of embedding models trained in one of the seven languages and tested on all seven. The results of this section demonstrate RISE multilingual performance computed by three embedding models. See Appendix B for comprehensive results across all phenomena for each model.
Negation emerges as the most robust discourse-level, semantic-syntactic transformation, achieving the highest mean rotor alignment score (0.788) across all model-language combinations with performance ranging from 0.686 to 0.918. Figure 2 demonstrates RISE performance on negation for each model. RISE transformations for negation are most geometrically consistent in text-embedding-3-large. Negation’s strong performance indicates that generalizable discourse-level, semantic-syntactic changes are captured by RISE and best applied cross-lingually in text-embedding-3-large.
@@
The cross-language analysis fully presented in Appendix B supports our hypothesis that discourse-level semantic-syntactic transformations correspond to geometric operations on the embedding manifold. The variation across models, preservation of linguistic relationships across languages, and transformation patterns indicate that RISE successfully identifies semantic-syntactic transformation on the embedding manifold. The limitations and future work are discussed further on.
Figure 3: Cross-language transfer heatmaps for text-embedding-3-large showing RISE performance across all language pairs for conditionality, negation, and politeness transformations. Darker colors indicate higher cosine similarity between predicted and target embeddings.
@@ 6.2 Cross-Model Transfer Comparison
To evaluate RISE prototypes’ robustness to transfer across different embedding architectures, we conducted cross-model mapping experiments using the method developed by Morris et al. (2020). This approach learns statistical mappings between embedding spaces through principal component analysis (PCA) and distributional alignment, enabling transfer of learned RISE prototypes from one model to another. We specifically examined transfer from text-embedding-3-large (3072-dimensional) to bge-m3 (1024-dimensional), demonstrating cross-model semantic transfer across different dimensionalities and training objectives. For each language pair and phenomenon, we learn RISE prototypes in text-embedding-3-large using 80% of the data, map these prototypes and $e_{1}$ to bge-m3 space, and evaluate performance on native bge-m3 embeddings using the remaining 20%. Figure 4 demonstrates comprehensive cross-model and cross-language transfer results.
@@ 6.3 English Task-Based Comparison
Our main investigation is how well RISE peforms in in multi-lingual settings. However there are limited external datasets for evaluating the performance discourse-level, semantic-syntactic transformation tasks. Due to the limited resources, we had to select the most related datasets, BLiMP and SICK. BLiMP is a paired sentence dataset for major grammatical phenomena in English, and SICK is a dataset with paired sentences with entailment, contradiction, and neutral labels.
Table 2 summarizes RISE performance across the three selected datasets. The results confirm that all models achieve strong performance, with particular strengths varying by dataset: mBERT excels on grammatical tasks (BLiMP) and contradiction detection (SICK), while bge-m3 shows the most consistent performance across synthetic multilingual data. The dramatic performance gap between BLiMP (>>0.92) and SICK (0.62-0.74) suggests that RISE rotations might be capturing something more specific than general semantic transformations.
The high BLiMP performance indicates RISE excels at preserving grammatical/syntactic structure, while the moderate SICK performance suggests these same rotations don’t preserve semantic relatedness as well. These results show that benchmark choice dramatically affects relative model ranking. Robustness depends on whether the task prioritizes cross-lingual consistency (favoring bge-m3) or raw performance on specific phenomena (favoring text-embedding-3-large for negation, mBERT for grammatical tasks).
Table 2: RISE Performance Across Three Datasets: The performance is measured with the rotor alignment score between RISE-steered embeddings and target embeddings where bold values indicate best performance per dataset. text-embedding-3-large is abbreviated as TE3L.
@@
6.4 Linear Baseline Comparisons
The full results presented in Appendix C compare RISE against two standard baselines, Mean Difference Vectors (MDV) and Procrustes alignment, across the same three datasets. MDV is not Euclidean. MDV preserves spherical structure and naturally resembles RISE more closely than Procrustes. This distinction is directly reflected in the results: MDV and RISE transfers best across languages where Procrustes fails.
The strongest performance appears in monolingual English evaluation (BLiMP), while performance drops substantially for Procrustes on semantic relatedness (SICK) shown in Table 3. This shift in performance reflects Procrustes’ inability to identify a generalizable semantic–syntactic relationship as expected by method. Procrustes fits a single global rotation which is too rigid for the cross-lingual and cross model analysis In contrast, RISE maintains stable cross-lingual and cross-model performance (e.g., App. B. Figures 5–7), indicating that geometric operations on the manifold better capture discourse-level semantic structure than Euclidean differences.
The MDV vs. RISE vs. Procrustes results reinforce our earlier claim that methods operating on the curved manifold (where sentence embeddings inherently reside) perform better than Euclidean/linear methods. Most steering and probing techniques operate in linear space, and we conjecture that this geometric mismatch helps explain why linear methods struggle to generalize. In short, Procrustes fits a single global rotation which is too rigid for the cross-lingual and cross model analysis. Geometric transformations, like RISE and MDV, are better suited for semantic-syntactic analysis and cross-lingual stability.
Method Monolingual Syntactic Monolingual Semantic Cross-Language Transfer
(BLiMP) (SICK) (All Phenomena)
RISE Strong (0.97) Strong (0.84) Moderate–Strong (0.74–0.89)
MDV Strong (0.97) Strong (0.83) Moderate–Strong (0.72–0.91)
Procrustes Strong (0.99) Moderate (0.67) Failing–Weak (0.25–0.62)
Table 3: Condensed summary of baseline comparisons from Appendix C using the cosine-similarity interpretation scale from Table 1. RISE and MDV show Strong monolingual and Moderate–Strong cross-language structure, whereas Procrustes drops to Weak or Failing consistency outside syntactic, same-language settings.
7 Discussion & Future Work
@@
Our findings demonstrate that meaningful semantic-syntactic operations can be recovered as geometric transformations in modern language model representations. RISE successfully identifies consistent geometric structure for discourse-level semantic-syntactic changes, primarily for text-embedding-3-large and negation in multilingual settings. The results demonstrating spherical methods, RISE and MDV, out perform linear methods, Procrustes alignment, provide positive results for extending the LRH to spherical spaces.
@@
Evaluation benchmarks (Table 2) reveal task-dependent effectiveness. RISE achieves near-perfect performance on syntactic acceptability (BLiMP: 0.93-0.96) but only moderate performance on semantic similarity (SICK: 0.62-0.74), suggesting better alignment with grammatical rather than semantic transformations. Section 6.1 shows that negation and conditionality are the most generalizable discourse-level, semantic-syntactic changes captured by RISE and best applied cross-lingually in text-embedding-3-large. Our cross-model transfer experiments expose an English-centric bias, with English achieving 20% higher transfer scores than languages like Zulu. This English-centric bias persists across all semantic transformations, indicating that current multilingual models encode geometric structures that prioritize English. Future work should focus on developing more equitable multilingual representations and investigating which language-specific geometric structures are an inherent feature of the models.
@@
Together these results support that RISE is most successful at identifying semantic transformation with distinct grammatical factors, but more work is needed to justify semantic transformations in multilingual models are universal geometric operations. First, our analysis focuses on three specific linguistic transformation types. Future work should expand to additional semantic and pragmatic phenomena to test the generality of geometric consistency principles. Second, while our experiments used three diverse embedding models (text-embedding-3-large, bge-m3, and mBERT), validation across additional architectures would strengthen claims about the universality of geometric semantic structure. Third, the reliance on GPT-4.5 for data generation may introduce subtle biases toward English-centric conceptualizations of semantic phenomena. Future work should incorporate more diverse data sources and validation by native speakers.
@@ 8 Conclusion
The ability to learn geometric transformations for discourse changes relates to work on text generation and steering vectors (Turner et al., 2023; Li et al., 2023). Our rotor-based approach, RISE, provides a geometric framework for understanding and improving interpretability in language models. This work investigated whether discourse-level semantic-syntactic transformations in multilingual embedding spaces correspond to intrinsic geometric operations, specifically rotations identified through the RISE method. Our comprehensive evaluation across multiple baselines, models, languages, and datasets reveals a more complex reality than initially hypothesized. This work demonstrates that modern language model representations maintain interpretable geometric structure for some semantic-syntactic transformations, extending the promise of geometric semantics from early word embeddings to contemporary transformer models. We show that:
@@
    Semantic transformations with clear syntactic mapping demonstrate the most consistent geometric structure.
@@
    RISE successfully identifies semantically meaningful geometric structure in high-dimensional embedding spaces that generalizes cross-lingually and across model architecture.
As language models continue to evolve, understanding these geometric foundations will be crucial for developing more interpretable AI systems. By revealing transferable geometric structure in semantic transformations (e.g. negation and conditionality), this work opens new possibilities for understanding language model behavior through geometric interventions. Our work promotes geometric methods as more appropriate approaches to cross-lingual semantic interpretation, achieving 77%-95% cross-language transfer effectiveness across typologically diverse languages. By developing RISE, we demonstrate that interpretable structure exists for some grammatically distinct semantic transformations, providing a tools for understanding how these systems encode semantic knowledge. While RISE remains valuable for analyzing model-specific semantic structures, claims about universal geometric operations require substantial qualification.
@@ Appendix B Cross-Language Transfer Analysis and Results
To test whether geometric transformations generalize across languages, we conducted comprehensive cross-language transfer experiments. This section reports detailed results across three models and three semantic phenomena, analyzing both quantitative performance and geometric properties of learned transformations.
@@ B.1 Cross-Language Transfer Performance
The above heatmaps demonstrate comprehensive cross-language transfer results across our three models. Training rotor prototypes on one language and evaluating on others reveals promising cross-linguistic performance, particularly for negation and conditionality. Most language pairs show transfer scores above 0.70, with negation achieving particularly strong off-diagonal performance (most scores >> 0.80).
@@ B.2 Geometric Analysis of Cross-Language Centroids
Analysis of the learned centroids reveals additional insights into the geometric structure of semantic transformations. For each phenomenon, we computed “ideal” transformation vectors by averaging canonicalized transformed embeddings across languages.
@@ B.3 Quantitative Cross-Language Analysis
Model Phenomenon All Transfers Monolingual Cross-Lang Ratio TE3L (3072d) Conditionality 13.6× ± 0.7 14.5× 13.5× 0.93 Negation 19.7× ± 1.2 20.6× 19.6× 0.95 Politeness 23.1× ± 1.9 25.3× 22.8× 0.90 bge-m3 (1024d) Conditionality 13.9× ± 0.6 14.5× 13.8× 0.95 Negation 18.5× ± 0.8 19.3× 18.4× 0.95 Politeness 25.2× ± 1.2 26.4× 25.1× 0.95 mBERT (768d) Conditionality 12.8× ± 1.3 15.0× 12.5× 0.83 Negation 18.0× ± 1.8 20.9× 17.5× 0.84 Politeness 20.8× ± 3.9 26.1× 20.0× 0.77
@@
Model Dims Validation Avg Cross-Lang Avg Random Adv TE3L 3072 0.774 19.0× 6.3× bge-m3 1024 0.790 19.8× 11.7× mBERT 768 0.802 16.9× 11.9×
@@
Tables 4 and 5 provide comprehensive quantitative analysis of cross-language transfer performance. Notably, all models maintain strong cross-language performance (77%–95% of monolingual performance), with bge-m3 showing the most consistent cross-language effectiveness across all phenomena.
Appendix C Linear Baselines Comparisons
This appendix reports the full results for the linear baseline comparisons requested by the reviewers. We thank the reviewers for this valuable suggestion as these results did strengthen our paper. We implemented two baselines: Procrustes alignment and Mean Difference Vectors (MDV). MDV is not truly Euclidean: it computes mean displacements using the manifold’s geometry (via log/exp maps), preserving spherical structure. Thus MDV functions naturally resembles RISE more closely than Procrustes. We evaluated them alongside RISE on three datasets: BLiMP, SICK, and our multilingual synthetic dataset.
The strongest performance appears in monolingual English evaluation (BLiMP), while performance drops substantially for Procrustes on semantic relatedness (SICK) shown in Table 3. This shift in performance reflects Procrustes’ inability to identify a generalizable semantic–syntactic relationship as expected by method. Procrustes fits a single global rotation which is too rigid for the cross-lingual and cross model analysis In contrast, RISE maintains stable cross-lingual and cross-model performance (e.g., App. B. Figures 5–7), indicating that geometric operations on the manifold better capture discourse-level semantic structure than Euclidean differences.
The MDV vs. RISE vs. Procrustes results reinforce our earlier claim that methods operating on the curved manifold (where sentence embeddings inherently reside) perform better than Euclidean/linear methods. Most steering and probing techniques operate in linear space, and we conjecture that this geometric mismatch helps explain why linear methods struggle to generalize. In short, Procrustes fits a single global rotation which is too rigid for the cross-lingual and cross model analysis. Geometric transformations, like RISE and MDV, are better suited for semantic-syntactic analysis and cross-lingual stability.
C.1 Cross-Language Transfer Heatmaps
Figures 810 show cross-language cosine similarity for the three semantic transformations (Conditionality, Negation, Politeness) under Mean Difference Vectors (MDV), Procrustes alignment, and RISE.
(a) MDV
(b) Procrustes
(c) RISE
Figure 8: Cross-language transfer for Conditionality across seven languages.
(a) MDV
(b) Procrustes
(c) RISE
Figure 9: Cross-language transfer for Negation across seven languages.
(a) MDV
(b) Procrustes
(c) RISE
Figure 10: Cross-language transfer for Politeness across seven languages.
C.2 Natural-Language Validation: BLiMP and SICK
Figure 11 reports mean cosine similarity on BLiMP (syntactic) and SICK (semantic) for the three methods.
Figure 11: Natural language validation on BLiMP (syntactic acceptability) and SICK (semantic relatedness) for RISE, MDV, and Procrustes. Error bars denote standard deviation across examples.
Appendix D Prompt Templates
@@
D.1 Neutral Sentence Generation
@@ D.4 Conditionality
Appendix E Data Generation Methodology
E.1 Diversity Controls
To guard against artifacts that might arise from narrow lexical or topical coverage, we apply several sampling diversity control. (1) Each neutral sentence prompt draws its situation description from a randomly chosen context category and exemplar, yielding a wide topical spread before any transformation is applied. (2) Within every language we shuffle sentence–feature assignments so that no specific lexical field correlates with a particular transformation subtype. (3) For each transformation we uniformly sample property values (e.g., negation particle, politeness strategy) per language and sentence, guaranteeing that every combination of language and subtype appears the same number of times. (4) After generation we remove near-duplicates and enforce a 5–25 token length window, which empirically yields a near-uniform length distribution. Together these steps ensure that our corpus varies in topic, syntax, and lexical choice while remaining balanced across languages and transformation subtypes. These controls ensure that observed geometric patterns reflect semantic properties rather than artifacts of lexical choice or sentence structure.